OpenStudy (anonymous):

Determine the equation of the function that has the indicated transformation applied to the function f(x)= 3(x+1)^2 -2

OpenStudy (anonymous):

a) translated up 2 and moved 2 units to the right

OpenStudy (anonymous):

a. f(x)= 3(x+1)^2-2 =3(x+1+2)^2-2+2 = 3(x-3)+0

OpenStudy (anonymous):

Can anybody tell me if this is correct?

OpenStudy (anonymous):

b) reflected in the x-axis and strtched by a factor of 3.

OpenStudy (anonymous):

b. f(x)= 3(x+1)^2-2 =-3(x+1+2)^2-2 = -6(x+2)^2-2

OpenStudy (anonymous):

3(x+1)^2-2 = 3x^2+6x+1= \[\frac{1}{3}\times3 \times (3x ^{2}+6x+1)= \frac{1}{3}\times ((3x)^{2}+18x+3) \] \[\frac{1}{3}\times((3x+3)^{2}-6) = \frac{1}{3}(3x+3)^{2} -2\]

OpenStudy (mertsj):

The vertex of the given function is (-1,-2) If you move it 2 right and 2 up the new vertex will be (1,0) and the equation will be: y= 3(x-1)^2+0

OpenStudy (mertsj):

If you apply a stretch factor of 3 and reflect in the x axis the new equation is: y = -9(x+1)^2-2